Chandler is 5 years older than Monica. Fifteen years ago, the sum of their
ages was 45. Find their present ages

Answer by ikleyn(52787)

.

If M is Monica's age, then Chandler's age is M+5 years.


Fifteen years ago


    (M-15) + ((M+5) - 15) = 45

     2M    -  25          = 45

     2M                   = 45 + 25 = 70

      M                             = 70/2 = 35.


ANSWER.  Monica is 35 years old; Chandler is 35+5 = 40 years old.